OpenStudy (happykiddo):

Trig substitution for int(9/(sqrt(3+x^2)))dx the answer is 9ln(x+sqrt(3+x^2))+C.

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{9}{\sqrt{3+x^2}}~dx}\)

OpenStudy (solomonzelman):

like this?

OpenStudy (happykiddo):

yeah

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle x=\sqrt{3}\cdot \tan \theta}\)

OpenStudy (happykiddo):

yeah I got that

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{9}{\sqrt{3+\left(\sqrt{3}\cdot \tan \theta\right)^2}}~\left(\sqrt{3}\cdot \sec^2\theta\right)d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{9\sqrt{3}\cdot \sec^2\theta}{\sqrt{3+3\tan ^2\theta}}~d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{9\sqrt{3}\cdot \sec^2\theta}{\sqrt{3}\sqrt{1+\tan ^2\theta}}~d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{9\sec^2\theta}{\sqrt{\sec ^2\theta}}~d\theta}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{9\sec^2\theta}{\sec \theta}~d\theta}\) \(\large\color{black}{\displaystyle9\int\limits_{~}^{~}\sec\theta~d\theta}\)

OpenStudy (solomonzelman):

I am not supposed to solve it for you, so try to work out the solution please. I will keep typing and I will store it somewhere

OpenStudy (solomonzelman):

You can't tell me the integral for sec \(\theta\) ?

OpenStudy (happykiddo):

ln (tan x + sec x)

OpenStudy (solomonzelman):

yes, correct

OpenStudy (happykiddo):

yeah I got it from here

OpenStudy (solomonzelman):

but in this case, NOT X, but THETA

OpenStudy (solomonzelman):

and +C

OpenStudy (happykiddo):

yeah theta

OpenStudy (solomonzelman):

Now, you have to solve for theta in terms of x, and substitute the original variable.

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle9\int\limits_{~}^{~}\sec\theta~d\theta= \\[0.9em] 9\ln\left(\tan \theta+\sec \theta\right)+C=\\[0.9em] 9\ln\left(\tan \left(?\right)+\sec \left(?\right)\right)+C}\) \(\large\color{blue}{\displaystyle x=\sqrt{3}\cdot \tan \theta}\) \(\large\color{blue}{\displaystyle x/\sqrt{3}=\tan \theta}\) \(\large\color{blue}{\displaystyle \tan^{-1}\left(x/\sqrt{3}\right)= \theta}\) \(\large\color{red}{\displaystyle9\ln\left(\tan \left(\tan^{-1}\left(x/\sqrt{3}\right)\right)+\sec \left(\tan^{-1}\left(x/\sqrt{3}\right)\right)\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(\frac{x}{\sqrt{3}}+\sec \left(\tan^{-1}\left(x/\sqrt{3}\right)\right)\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(\frac{x}{\sqrt{3}}+\sqrt{1+\arctan^2 \left(\tan^{-1}\left(x/\sqrt{3}\right)\right)}\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(\frac{x}{\sqrt{3}}+\sqrt{1+\left(\frac{x}{\sqrt{3}}\right)^2}\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(\frac{x}{\sqrt{3}}+\sqrt{1+\frac{x^2}{3}}\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(\frac{x}{\sqrt{3}}+\sqrt{\frac{x^2+3}{3}}\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(\frac{x}{\sqrt{3}}+\frac{\sqrt{x^2+3}}{\sqrt{3}}\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(\frac{x+\sqrt{x^2+3}}{\sqrt{3}}\right)+C}\) \(\large\color{black}{\displaystyle9\ln\left(x+\sqrt{x^2+3}\right)-9\ln\left(\sqrt{3}\right)+C}\) 9ln(√3) is also a constant, and an arbitrary constant minus 9ln(√3), will still yield an arbitrary constant. THereofre you have: \(\large\color{black}{\displaystyle9\ln\left(x+\sqrt{x^2+3}\right)+C}\)

OpenStudy (happykiddo):

Yeah I was making sec theta into 1/cos theta. Then integrating thanks

OpenStudy (solomonzelman):

yes, that is certainly the way to show the integral of secθ, knowing the fact that: sec θ = cos θ / (1 - sin\(^2\)θ)

OpenStudy (solomonzelman):

then u=sin θ and so on...